OpenStudy (anonymous):

Can yall help me out with this 8. p to p' (2, -6) for the glide reflection where the translation is (x, y) to (x+2, y-2) and the line of reflection is y= -x. what are the coordinates of P? (11, -3) (4, 0) (-11,3) (-4, 0)

OpenStudy (anonymous):

do you have a graph?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

um well lets see

OpenStudy (anonymous):

here ill get you a gragh

OpenStudy (anonymous):

try it with this

OpenStudy (anonymous):

@Bghostrider2

OpenStudy (anonymous):

sorry chrome crashed

OpenStudy (anonymous):

and ok

OpenStudy (anonymous):

dude i have know idea how to do the problem

OpenStudy (anonymous):

hold on gonna switch to explorer

OpenStudy (anonymous):

okay well um give me just 1 minute and ill get back to you

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

alright im back

OpenStudy (anonymous):

okay so first try and graph the problem on the graph paper

OpenStudy (anonymous):

my main problem is the y=-x...I don't know hw to change the coordinates with it

OpenStudy (anonymous):

when you do go ahead and tell me what point on the graph reflects (2,-6)

OpenStudy (anonymous):

no put all the points on the graph and then see what one reflects that problem

OpenStudy (anonymous):

there are only 2 points

OpenStudy (anonymous):

even A,B,C,and D put those on the graph as well

OpenStudy (anonymous):

ok I put them on the graph

OpenStudy (anonymous):

now what one reflects the point (2,-6)

OpenStudy (anonymous):

none

OpenStudy (anonymous):

okay then what i would do is put the points in the equation and see what you get

OpenStudy (anonymous):

okay well im also sort of confused lets see if we can find some one that will know how to set up the equation

OpenStudy (anonymous):

would it be option A

OpenStudy (anonymous):

uh yes actualy i just set it up and i got a also

OpenStudy (anonymous):

well....it wasn't a

OpenStudy (freckles):

so we started at a point (x,y) and we want to end up at (2,-6) Well we can go backwards from that point back through operations so let's back through the reflection over y=-x that is if you have (x,y) then you will have (-y,-x) if you reflect it over that line. So when you reflect (2,-6) about y=-x what point do you get? Once you find that point we have to take back through the translation to get where you started at.

OpenStudy (freckles):

end point (2,-6) after reflection through y=-x we have (6,-2) I will do the reflection part for you now we need to figure out what is x and what is y when (6,-2)=(x+2,y-2) you just have two equations to solve now

OpenStudy (anonymous):

it was b

OpenStudy (freckles):

exactly right good job

OpenStudy (freckles):

since 6=x+2 gives x=4 and -2=y-2 gives y=0